OpenStudy (anonymous):

Show that that two lines x-2y+7=0 and 2x+y-16=0 are perpendicular each other need Working out please

OpenStudy (anonymous):

you can use the fact that product of slopes of two perpendicular line is -1.

OpenStudy (anonymous):

well what is it in slope intercept form??

OpenStudy (anonymous):

just write it out for me

OpenStudy (anonymous):

can you reduce the equations above to y=mx+c form? If yes, the m part is slope.

OpenStudy (anonymous):

are you still confused because i can show you step by step

OpenStudy (anonymous):

yea im pretty much confused

OpenStudy (anonymous):

ok if you want to change a equation into a slope intercept form you must leave the y alone like this equation (x-2y+7=0) you must leave the y alone (x - x - 2y + 7 - 7 = 0 - x -7) which becomes (2y = -x - 7) =(y = -1/2(x) - 7/2)

OpenStudy (anonymous):

try the second equation

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

yea im here im doing consumer arithmetics

OpenStudy (anonymous):

hmmm...